### v2

@@ Abstract.
Subgraph isomorphism is a fundamental problem in graph analysis that seeks to find all instances of a pattern graph within a larger data graph while preserving structural relationships. This NP-complete problem is essential across diverse computational domains, from biological network analysis and social network mining to quantum circuit optimization. Traditional approaches rely on backtracking algorithms such as VF2, which suffer from inherent sequential bottlenecks that limit their ability to leverage modern parallel computing architectures. In this work, we introduce $\Delta$-Motif, a novel GPU-accelerated subgraph isomorphism algorithm that reformulates the problem through the lens of database operations. Our key insight is in representing both data and pattern graphs in tabular format, transforming subgraph isomorphism into fundamental database primitives including joins, sorts, merges, and filters. $\Delta$-Motif decomposes graphs into small building blocks called motifs and systematically combines these using scalable relational operations. By leveraging mature, highly optimized libraries from the NVIDIA RAPIDS ecosystem and Pandas framework, our solution achieves massive parallelization while remaining portable across any database system that supports standard relational primitives. Through benchmarking, we show that $\Delta$-Motif significantly outperforms established algorithms like VF2, achieving speedups of up to $595\times$ on GPU architectures. We demonstrate the practical impact of this approach by applying it to quantum circuit compilation, addressing a key bottleneck in quantum computing workloads and unlocking scaling to devices in the near and medium term. Our approach democratizes high-performance graph processing by making it accessible through familiar database abstractions, eliminating the need for specialized low-level programming while delivering exceptional computational efficiency.
PVLDB Reference Format:
PVLDB, 14(1): XXX-XXX, 2020.
doi:$X$.XX/XXX.XX This work is licensed under the Creative Commons BY-NC-ND 4.0 International License. Visit https://creativecommons.org/licenses/by-nc-nd/4.0/ to view a copy of this license. For any use beyond those covered by this license, obtain permission by emailing info@vldb.org. Copyright is held by the owner/author(s). Publication rights licensed to the VLDB Endowment.
Proceedings of the VLDB Endowment, Vol. 14, No. 1 ISSN 2150-8097.
doi:$X$.XX/XXX.XX
@@ 1. Introduction
Graphs have emerged as one of the most flexible and powerful data structures to represent complex relationships, leading to their ubiquitous presence across various computational domains. From modeling molecular interactions in bioinformatics (BioinformaticsAndGraphs, 1) and tracking social connections in network analysis (SocialNetworksAndGraphs, 2) to mapping dependencies in cybersecurity systems (CybersecAndGraphs, 3), graphs provide an intuitive and mathematically rigorous framework for capturing the intricate patterns that define real-world systems.
@@
Central to many graph-based applications is the subgraph isomorphism problem, which seeks to find all instances of a pattern graph within a larger data graph while preserving structural relationships. This problem is known to be NP-complete (Cook1971, 4), yet it remains essential for tasks ranging from motif discovery in biological networks (BiochemAndSubgraphIso, 5), social networks analysis (Dasgupta2020DiscoveringIS, 6), data mining (Zhao2012LargeSC, 7), drug discovery(Aung2008BSAlignAR, 8) and circuit mapping in quantum devices (Nation2023, 9). The ubiquity of its application to various use cases can create critical bottlenecks in compute and workload execution. Driving efficiencies for subgraph isomorphism will alleviate key barriers to the incorporation of emerging technologies like quantum computing into high-performance computing environments, and thereby its adoption for meaningful impact.
Traditional approaches to subgraph isomorphism have relied on backtracking algorithms such as VF2 (Cordella2004, 10), VF2++ (Juttner2018, 11) and VF3 (Carletti2019VF3L, 12). These approaches rely on an almost DFS-like traversal pattern, searching for subgraphs, which makes most of this solution inherently sequential. Code specialization techniques have arisen as well (Dryadic2021, 13) to improve performance.
A key use case we are addressing in this work is the application of graph algorithms in the domain of quantum circuit compilation. Recent work by Nation et al. (Nation2023, 9) demonstrated the use of a Rust-based implementation of VF2 (Rustworkx2022, 14) in such a domain, highlighting both the utility and limitations of these classical approaches. VF2 belongs to a family of algorithms that rely fundamentally on backtracking search strategies to explore the solution space. As we explored VF2 in our solution, we found backtracking-based algorithms to be challenging to parallelize effectively, which limited our ability to use hundreds of CPU cores and current GPU architectures.
@@
To overcome the sequential bottlenecks and limited parallelism of existing subgraph isomorphism enumeration, we developed a novel data-science centric solution that reformulates of the problem as a series of tabular data operations. By representing both data and pattern graphs in tabular format, subgraph isomorphism is transformed into fundamental database operations such as joins, sorts, merges and filters. This reformulation enables us to leverage mature, highly-optimized open-source and commodity libraries from the data science and database communities, transforming a traditionally graph-theoretic problem into one that naturally benefits from established data processing ecosystems.
@@
Building on this tabular reformulation, another key strength is its portability across tabular data systems. The algorithm can be implemented on any database platform that supports these primitives, from traditional relational databases to in-memory tabular representations such as Pandas (mckinney2011pandas, 15). As the result, it allows us to seamlessly leverage the NVIDIA RAPIDS ecosystem  (Cupy2017, 16, 17, 18) for GPU acceleration, while maintaining full compatibility with the popular Pandas (mckinney2011pandas, 15) data science framework when GPUs are not readily available.
Although the aforementioned operations are conceptually straightforward and also fundamental to database systems, achieving high performance in practice is nontrivial. For this reason, we chose to build upon RAPIDS as a well-tuned library, ensuring that future performance improvements to the library will directly benefit our algorithm’s performance. By leveraging these primitives in our design, we achieve massive parallelism and deliver strong performance on NVIDIA GPUs. To the best of our knowledge, this is the first implementation of such an approach to solve the subgraph isomorphism problem, demonstrating how fundamental database operations can serve as a powerful foundation for solving computational tasks well beyond their traditional scope.
@@ 1.1. Contributions
We organize our contributions into four key areas. First, we propose a novel algorithm for subgraph isomorphism enumeration based on motif-driven decomposition. Instead of directly matching full subgraphs, our algorithm constructs solutions by decomposing graphs into reusable building blocks we call motifs: small, well-defined subgraph patterns such as paths or triangles. By systematically combining these indivisible units, the algorithm reconstructs complete matches in a modular and scalable manner. As it will become apparent in a following section, this is the basis for the name of our new algorithm: $\Delta$-Motif.
[…]
Third, we demonstrate the impact of the $\Delta$-Motif by applying it to quantum circuit compilation, where current superconducting quantum processors require efficient mapping between the circuit and the device’s coupling architecture. Our technique enables circuit compilation on significantly larger devices compared to what is possible with current methods, enabling the scalability needed for next-generation quantum processors.
Fourth, we show that the $\Delta$-Motif algorithm achieves high performance, both on CPU and GPU, while using commodity and open-source software. Our high-performance solution does not require extensive low-level hardware knowledge or manually tuning any high-performance code.
By reformulating the subgraph isomorphism problem in terms of tabular data and relational operations, the algorithm inherits massive parallelism from existing optimized libraries. This approach democratizes high-performance graph processing by making it accessible through familiar data science tools and abstractions, while still delivering the computational efficiency brought by large-scale parallelism.
@@ 2. Related Work
The VF2 algorithm (Cordella2004, 10) uses a depth-first backtracking strategy to incrementally build partial isomorphisms between graphs, applying pruning techniques at each step to maintain consistency. It is specifically designed for improved memory efficiency and performs well on dense graphs. Similar backtracking-based approaches have been exploited by specialized algorithms such as AutoMine (AutoMine2019, 19), which generates optimized code targeting specific graph patterns and Dryadic (Dryadic2021, 13), which uses compilation techniques to further generalize on the idea of specialized code per query.
@@
While VF2 has demonstrated utility on relatively small graphs, its performance degrades significantly as both the size and density of the target graph increase. Several variants have attempted to address these limitations: VF2++ (Juttner2018, 11), VF3 (Carletti2017, 20) and VF3L (Carletti2019VF3L, 12) improve search efficiency through more aggressive pruning techniques, while VF3P (Carletti2019VF3P, 21) extends these efforts by adding CPU-based parallelization. Although these improvements help accelerate the discovery of individual solutions, they do not fundamentally address the limitations of the tree-based search structure. More critically, as hardware development has increasingly focused on architectures that support massive parallelism, these backtracking-based designs do not map well to modern GPU architectures due to their inherently sequential nature and limited parallelization opportunities.
While recent GPU-based approaches (FastSubgraphmatchingGPU2015, 22, 23) have achieved performance improvements through specialized strategies like breadth-first exploration, vertex oriented strategies, and specialized data structures, they require custom GPU kernels, hardware-specific optimizations, and extensive low-level tuning that creates significant barriers to adaptation. In contrast, our approach prioritizes programmability and long-term maintainability over short-term performance gains.
Our work also draws ideas from a different subfield in graph processing. The concept of motifs is well established in graph analytics, where motifs have been applied to understand the recurrent structures in biological networks(MotifsAndBio, 24) and network classification tasks(MotifsAndClassif, 25). However, most existing work has focused on motif discovery and identification rather than leveraging motifs as computational building blocks to solve graph-centric problems such as subgraph isomorphism.
The graph database community has also explored this domain through indexing approaches (FG-Index-Graph-Databases, 26, 27), these efforts focus on building indices of frequent subgraphs in a database to accelerate subgraph isomorphism queries. Their key insight is that decomposing subgraph isomorphism into two phases: indexing and verification, can significantly accelerate query response times. Our work adopts a conceptually similar two-phase approach; however, we apply this decomposition strategy to exploit structural similarities within individual graphs.
@@ 2.1. Subgraph Isomorphism Enumeration in Quantum Circuit Compilation
As quantum hardware steadily advances, near term processors have emerged as promising platforms for exploring algorithmic realizations to demonstrate quantum advantage (Preskill2018, 28, 29, 30). However, these systems remain fundamentally constrained by imperfect control (Krantz2019, 31, 32, 33, 34), hardware variability (Ravi2023, 35, 36), and short coherence times (Stilck2021, 37). Quantum circuit compilation, plays a critical role in minimizing operational overhead and mitigating error accumulation to improve execution fidelity (Nation2025, 38). Given a quantum device and a quantum circuit, the compilation pipeline aims to realize the target algorithm with fewer operations and a shorter execution time, thereby reducing the impact of noise. These procedures collectively define the topology and operation count of the compiled circuit before execution on hardware.
@@
Crucially, not all qubits (the information carriers in quantum computers) in a quantum processor offer the same operational quality (Krantz2019, 31, 36). Due to the inherent fragility of quantum systems and the challenges associated with device calibration, processors often contain faulty or low-fidelity qubits and couplings. As a result, the choice of which subset of physical qubits to utilize for executing a compiled circuit can significantly influence the circuit’s overall computational fidelity. For the same circuit, different mappings of circuit logical qubits into hardware physical qubits, referred to as layouts, may yield significantly different execution outcomes (Nation2023, 9, 39). This sensitivity to choice of layout has motivated the development of systematic layout selection methods, typically structured as a two-phase process: layout generation and score-based ranking. In the layout generation phase, the search space of candidate mappings is explored by identifying subgraph isomorphisms between the circuit’s interaction graph and the hardware’s connectivity graph. In the ranking phase, each valid layout is evaluated using device-specific quality metrics to select the most promising layout. This two-phase approach enables the efficient identification of hardware-aware circuit configurations optimized for execution fidelity.
In addition to graph-theoretic methods, machine learning approaches have been explored to support layout selection. For example, Hartnett et al. (Hartnett2024, 39) introduced a learning-based model that ranks candidate layouts based on device-specific calibration data. However, such approach does not address the most challenging aspect of the problem, layout generation, which is the focus of our work. Instead, it relies on existing algorithms, such as VF2, to enumerate the full set of candidate layouts. The model is limited to the ranking of these precomputed options, and its effectiveness depends on large volumes of hardware-specific training data, which are expensive and time-consuming to obtain. Furthermore, the trained models are hardware-specific and do not generalize to different devices, limiting their broader applicability.
In this work, we present an alternative implementation of one of the most time-consuming stages in a state-of-the-art quantum circuit compiler. Our goal is to accelerate the layout selection task by applying a data-centric subgraph isomorphism algorithm to a problem traditionally approached from a graph-theoretic perspective. We further extend this algorithm to support highly efficient layout scoring and ranking capabilities, enabling an end-to-end demonstration of our approach on realistic quantum compilation tasks.

### v3

@@ Abstract.
Subgraph isomorphism is a fundamental problem in graph analysis that seeks to enumerate all occurrences of a pattern graph within a larger data graph while preserving structural relationships. This NP-complete problem underpins a wide range of applications, from biological and social network analysis to quantum circuit compilation. Classical algorithms such as VF2 rely on backtracking-based search, which suffers from inherent sequential bottlenecks and limits scalability on modern parallel hardware. In this work, we introduce $\Delta$-Motif, a novel GPU-accelerated subgraph isomorphism algorithm that reformulates the problem through the lens of database operations. Our key insight is in representing both data and pattern graphs in tabular format, transforming subgraph isomorphism into fundamental database primitives including joins, sorts, merges, and filters. $\Delta$-Motif decomposes graphs into small building blocks called motifs and systematically combines these using scalable relational operations. By leveraging mature, highly optimized libraries from the NVIDIA RAPIDS ecosystem and Pandas framework, our solution achieves massive parallelization while remaining portable across any database system that supports standard relational primitives. We evaluate $\Delta$-Motif on real-world network datasets and graph workloads motivated by quantum computing, which capture the structural properties and scales of subgraph isomorphism problems arising in quantum circuit compilation. Through benchmarking, we show that $\Delta$-Motif substantially outperforms established methods such as VF2 and GSI, achieving speedups of up to $595\times$ on GPU architectures. By eliminating the need for specialized low-level kernels, our data-centric formulation democratizes high-performance subgraph isomorphism, making scalable graph analytics accessible through familiar database abstractions.
@@ 1. Introduction
Graphs have emerged as one of the most flexible and powerful data structures to represent complex relationships, leading to their ubiquitous presence across various computational domains. From modeling molecular interactions in bioinformatics (undef, ) and tracking social connections in network analysis (undefa, ) to mapping dependencies in cybersecurity systems (undefb, ), graphs provide an intuitive and mathematically rigorous framework for capturing the intricate patterns that define real-world systems.
@@
Central to many graph-based applications is the subgraph isomorphism problem, which seeks to find all instances of a pattern graph within a larger data graph while preserving structural relationships. This problem is known to be NP-complete (undefc, ), yet it remains essential for tasks ranging from motif discovery in biological networks (undefd, ), social networks analysis (undefe, ), data mining (undeff, ), drug discovery(undefg, ) and circuit mapping in quantum devices (undefh, ). The ubiquity of its application to various use cases can create critical bottlenecks in compute and workload execution. Driving efficiencies for subgraph isomorphism will alleviate key barriers to the incorporation of emerging technologies like quantum computing into high-performance computing environments, and thereby its adoption for meaningful impact.
Traditional approaches to subgraph isomorphism, such as VF2 (undefi, ), VF2++ (undefj, ), and VF3 (undefk, ), rely on backtracking algorithms with a largely DFS-like traversal pattern. This search strategy makes these methods inherently sequential and difficult to parallelize effectively. Code specialization techniques have been proposed to mitigate these limitations (undefl, ), and a range of systems have explored distributed execution, accelerator-based designs, and commercial data-processing frameworks such as Hadoop (undefm, ), Cassandra (undefn, ), and Spark (undefo, ). Subgraph isomorphism workloads also vary widely in structure, including weighted or unweighted graphs, small structured patterns, large random patterns, and data graphs with diverse connectivity properties. As a result, no single algorithm is optimal across all settings, and existing work spans a broad set of complementary approaches. We discuss these trade-offs in more detail in Sec. 2.
A key motivation for this work is the use of subgraph isomorphism in quantum circuit compilation. Recent work by Nation et al. (undefh, ) applied a Rust-based implementation of VF2 (undefp, ) in this setting, illustrating both the practicality and limitations of classical backtracking approaches. In our own evaluation, we found that such algorithms limit effective utilization of hundreds of CPU cores and modern GPU architectures. In quantum computing, the pattern graphs of interest are relatively large, typically on the order of dozens to around one hundred vertices in state-of-the-art workloads. As devices continue to scale, even larger patterns are becoming relevant. At these sizes, many existing parallel subgraph isomorphism algorithms perform poorly or fail to scale effectively, whereas our algorithm remains efficient.
@@
To overcome the sequential bottlenecks and limited parallelism of existing subgraph isomorphism enumeration, we developed a novel data-science centric solution that reformulates of the problem as a series of tabular data operations. By representing both data and pattern graphs in tabular format, subgraph isomorphism is transformed into fundamental database operations such as joins, sorts, merges and filters. This formulation transforms a traditionally graph-theoretic problem into one that naturally aligns with modern data-processing ecosystems.
@@
A key strength of this approach is its portability across tabular data systems. The algorithm can be implemented on any database platform supporting standard relational operations, ranging from traditional databases to in-memory frameworks such as Pandas (undefq, ). This portability enables seamless acceleration using the NVIDIA RAPIDS ecosystem  (undefr, ; undefs, ; undeft, ) for GPU acceleration, while maintaining full compatibility with CPU-based execution.
Although the aforementioned operations are conceptually straightforward and also fundamental to database systems, achieving high performance in practice is nontrivial. For this reason, we chose to build upon the RAPIDS library, ensuring that future performance improvements to the library will directly benefit our algorithm’s performance while remaining deployable on commodity hardware using open-source software. As a result, the algorithm inherits massive parallelism from existing optimized libraries, delivering strong GPU performance and making high-performance graph processing accessible through familiar data science tools and abstractions, without requiring specialized low-level programming.
@@ 1.1. Contributions
We organize our contributions into four key areas. First, we introduce a novel algorithm for subgraph isomorphism enumeration based on motif-driven decomposition. Instead of matching full subgraphs directly, $\Delta$-Motif decomposes graphs into reusable building blocks, called motifs, such as paths and small cycles, and reconstructs complete matches by systematically combining these units using relational operations. While prior work has explored motif-based decomposition, our approach differs fundamentally by formulating subgraph isomorphism as a sequence of join-and-filter operations over tabular data.
Second, through extensive benchmarking, we show that $\Delta$-Motif significantly outperforms well-established algorithms such as VF2 and state-of-the-art GPU implementations including GSI across a wide range of data and pattern graph sizes. On GPU architectures, $\Delta$-Motif achieves speedups of up to $595\times$, with particularly strong performance on regular graph topologies with predictable connectivity, including those motivated by quantum-hardware architectures. We further demonstrate that careful motif selection can substantially improve performance, yielding up to $10\times$ speedups over naive strategies.
Third, we evaluate $\Delta$-Motif on two complementary benchmark suites. We first study real-world network datasets using small pattern graphs, such as triangle enumeration, which represent common graph analytics workloads. We then conduct extensive experiments on graph workloads motivated by quantum computing, where large pattern graphs with dozens to hundreds of vertices must be matched against structured data graphs. Together, these benchmarks demonstrate that $\Delta$-Motif performs robustly across both small-pattern and large-pattern regimes, achieving strong and consistent performance improvements over existing methods.
Fourth, we that $\Delta$-Motif achieves high performance on both CPU and GPU using commodity hardware and open-source software. By relying exclusively on standard data science and database abstractions, our algorithm does not require custom kernels and extensive low-level hardware tuning as compared to many existing GPU-based solutions (discussed in more detail in Sec. 2).
@@ 2. Related Work
The VF2 algorithm (undefi, ) uses a depth-first backtracking strategy to incrementally construct partial isomorphisms while applying pruning rules to maintain consistency. It is specifically designed for memory efficiency and performs well on relatively small or dense graphs. Similar specialized approaches include AutoMine (undefu, ), which generates optimized code for specific graph patterns, and Dryadic (undefl, ), which uses compilation-based techniques to accelerate pattern-specific subgraph matching. More broadly, a large body of CPU-oriented work explores pruning strategies, memory reduction, and algorithmic refinements (undefv, ; undefw, ; undefx, ; undefy, ). While effective in many settings, these approaches offer limited opportunities for scalable parallelism.
@@
Although VF2 performs well on relatively small graphs, its performance degrades as the size and density of the target graph increase. Several variants have attempted to address these limitations: VF2++ (undefj, ), VF3 (undefz, ), and VF3L (undefk, ) improve search efficiency through more aggressive pruning strategies, while VF3P (undefaa, ) introduces CPU-based parallelization. VF2-PS (undefab, ) extends the VF2 algorithm to work in Arachne (undefac, ) and using the Chapel programming language. .Although these extensions can accelerate the discovery of individual solutions, they do not fundamentally address the limitations of the tree-based search structure. As a result, these approaches offer limited scalability on modern hardware architectures designed to exploit massive parallelism, particularly GPUs.
Recent work has explored GPU acceleration for subgraph isomorphism (undefad, ; undefae, ; undefaf, ; undefag, ; undefah, ; undefai, ; undefae, ; undefaj, ; undefak, ). These methods achieve performance improvements through techniques such as breadth-first expansion, vertex-centric execution, and specialized data structures. However, most rely on custom GPU kernels, hardware-specific optimizations, and extensive low-level tuning that creates significant barriers to adoption. GAMMA (undefaj, ), for example, uses an out-of-core design spanning host and device memory to support larger graphs, but incurs performance penalties due to frequent data movement over the PCIe interconnect. In contrast, Decomine (undefal, ) uses a compiler-based approach to generate pattern-specific code, avoiding custom kernels at the cost of reduced generality.
Complementary to accelerator-based approaches, the graph database community has explored scalability through indexing and multi-phase execution strategies. Indexing-based techniques (undefam, ; undefan, ) accelerate queries by decomposing subgraph isomorphism into indexing and verification phases, while BICE (undefao, ) separates preprocessing from enumeration to reduce redundant computation. Our algorithm also adopts a multi-phase structure, yet it differs in its data-centric formulation and execution model.
A complementary direction focuses on scalability through distributed and disk-based execution models (undefap, ; undefaq, ; undefn, ; undefm, ; undefo, ), often using bulk synchronous processing. While such systems can scale to large datasets, they typically incur higher runtime overheads due to I/O and communication costs, and therefore commonly target smaller pattern sizes.
Beyond these system-level approaches, our work also draws ideas from a different subfield in graph processing. The concept of motifs is well established in graph analytics to characterize recurrent structures in biological networks(undefar, ) and network classification tasks(undefas, ). However, prior work has primarily focused on motif discovery and counting, rather than using motifs as computational building blocks for solving subgraph isomorphism. EGSM (undefat, ) and ESCAPE (undefau, ) decompose matching tasks into smaller components, an idea related in spirit to our approach. However, these methods rely on specialized data structures and search strategies, whereas our algorithm formulates decomposition and recombination entirely through relational operations.
Several existing approaches employ join-based formulations, including DGSM (undefag, ) and FGSM (undefai, ). Lai et al. (undefav, ) propose a multi-join strategy that reduces intermediate results by filtering rows between joins. Although related in spirit, their approach assumes join-only execution, whereas our method introduces additional synchronization and filtering steps between joins to enforce structural constraints. Many GPU-based methods, including DGSM (undefag, ), EGSM (undefat, ), and cuTS (undefaf, ), further rely on specialized data structures to store intermediate results, increasing implementation complexity and reducing portability across evolving hardware platforms. In contrast, our algorithm represents both intermediate and final results using standard tabular data structures, benefiting directly from ongoing optimizations in data processing libraries while remaining accessible to practitioners familiar with standard data science tools. This data-centric formulation also naturally extends to distributed environments through existing frameworks that manage partitioning and communication without requiring specialized synchronization logic.
Finally, existing solutions target a wide range of pattern types, differing in size, structure, and associated properties, which makes direct comparisons difficult. For instance, many studies focus on enumerating relatively small patterns, often fewer than a dozen vertices, even when operating on large data graphs. For example, EGSM (undefat, ) reports results for patterns of up to 16 vertices, while cuTS (undefaf, ) and ESCAPE (undefau, ) evaluate patterns of up to 5 vertices, and other work emphasizes small clique enumeration (undefah, ). At the same time, algorithm performance is strongly influenced by properties of the data graph itself. For instance, cuTS (undefaf, ) targets sparse networks such as road graphs, whereas SGSI (undefak, ) operates on weighted vertices and edges, introducing different computational trade-offs compared to unweighted settings.
These differences highlight the need for a range of approaches tailored to different application requirements and usage patterns. While our algorithm might not be the most performant in all cases, it shows strong performance in many cases (including outperforming many of the state-of-the-art solutions). One of the big benefits of new algorithm is that it avoids custom kernels and specialized data structures by relying on open-source tabular data frameworks with well-defined APIs, prioritizing programmability, portability, and long-term maintainability while maintaining strong performance on modern parallel hardware.
@@ 2.1. Subgraph Isomorphism Enumeration in Quantum Circuit Compilation
As quantum hardware steadily advances, near-term processors have emerged as promising platforms for exploring algorithmic realizations to demonstrate quantum advantage (undefaw, ; undefax, ; undefay, ). However, these systems remain fundamentally constrained by imperfect control (undefaz, ; undefaaa, ; undefaab, ; undefaac, ), hardware variability (undefaad, ; undefaae, ), and short coherence times (undefaaf, ). Quantum circuit compilation plays a critical role in minimizing operational overhead and mitigating error accumulation to improve execution fidelity (undefaag, ). Given a quantum device and a quantum circuit, the compilation pipeline aims to realize the target algorithm with fewer operations and a shorter execution time, thereby reducing the impact of noise. These procedures collectively define the topology and operation count of the compiled circuit before execution on hardware.
@@
Crucially, not all qubits (the information carriers in quantum computers) in a quantum processor offer the same operational quality (undefaz, ; undefaae, ). Due to the inherent fragility of quantum systems and the challenges associated with device calibration, processors often contain faulty or low-fidelity qubits and couplings. As a result, the choice of which subset of physical qubits to utilize for executing a compiled circuit can significantly influence the circuit’s overall computational fidelity. For the same circuit, different mappings of circuit logical qubits into hardware physical qubits, referred to as layouts, may yield significantly different execution outcomes (undefh, ; undefaah, ). This sensitivity to choice of layout has motivated the development of systematic layout selection methods, typically structured as a two-phase process: layout generation and score-based ranking. Layout generation explores search space of candidate mappings by enumerating subgraph isomorphisms between the circuit’s interaction graph and the hardware’s connectivity graph, after which valid layouts are evaluated using device-specific quality metrics to select the most promising one.
In addition to graph-theoretic methods, machine learning approaches have also been explored for layout selection. For example, Hartnett et al. (undefaah, ) proposed a learning-based model that ranks candidate layouts using device-specific calibration data. However, such approaches operate on precomputed candidate layouts and therefore do not address layout generation, the most computationally demanding stage of the problem.
In this work, we present an algorithm based on tabular operations motivated by graph workloads arising in quantum circuit compilation, where layout generation is formulated as a subgraph isomorphism enumeration problem. A distinctive characteristic of this domain is that the pattern graphs are relatively large, typically on the order of dozens to around one hundred vertices in state-of-the-art workloads. This contrasts with many traditional subgraph matching workloads that emphasize small pattern graphs (with 10 or fewer vertices) embedded within very large data graphs; however, large pattern graphs can still induce substantial search spaces and many valid embeddings, making enumeration computationally challenging. As quantum devices continue to scale and incorporate additional qubits, even larger interaction patterns are expected to become relevant, motivating algorithms capable of efficiently handling larger pattern graphs while exploiting the structured connectivity commonly observed in quantum hardware topologies.